OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2?

OpenStudy (aroub):

y1/x1=y2/x2 8/-6=y/-2 8(-2)=-6(y) -16=-6y y=8/3 or y=kx 8=k (-6) k==-4/3 y=kx y=-4/3 ( -2) y=8/3